OpenStudy (anonymous):

Select the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form. A. (y + 1) = 3(x – 2) B. (y – 1) = 3(x + 2) C. (x + 2) = 3(y – 1) D. (x – 2) = 3(y + 1)

OpenStudy (anonymous):

point slope form : y - y1 = m(x - x1) slope(m) = 3 (-2,1) x1 = -2 and y1 = 1 now we sub y - 1 = 3(x - (-2) y - 1 = 3(x + 2)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

its quite alright....glad I could help :)